# H line shape and radius-luminosity relation in 2.5D FRADO

### v1

@@ Abstract
Context. Galaxies with active galactic nuclei (AGN) exhibit broad emission lines as a key spectral feature. The shape of emission-line profiles depends on the complex dynamics of discrete clouds within a spatially extended region known as the Broad Line Region (BLR). The distribution of cloud positions within BLR, or the geometry of BLR indeed, is directly linked to measurements of time lags of BLR.
Aims. In this paper, we convolve a large grid of physically-based simulations of cloud distributions in BLR with photon-flux weighted emissivity of BLR clouds to investigate the generic shape of spectral line profiles. More importantly, we extract the time-delay histograms of corresponding models to calculate the size of BLR.
Methods. Our physical model is based on the assumption that the clouds are launched by the radiation pressure acting on dust in the atmosphere of the outer disk. It has very few global parameters. The model is appropriate for the low ionization part of the BLR, as it was shown by earlier model tests. It uses a non-hydrodynamical single-cloud approach to the BLR dynamics. In this way we simulate the distribution of positions and velocities of the clouds.
Results. We found that the width of line profiles gets broader with black hole mass, or with viewing angle, and gets narrower with accretion rate. The blue wing of the emission line profiles becomes more pronounced with increasing black hole mass and accretion rate, consistent with the formation and intensification of an outflow structure. We also found that the peak time-delays rather than averaged delay values better represents the observational trend and also the scatter in the radius-luminosity relation.
@@
Active Galaxies – Broad Line Region – Main Sequence of Quasars – time-delays – Eigenvector 1 – Broad Emission Lines1 Introduction
@@
Broad emission lines in the optical-UV spectra of active galactic nuclei (AGNs) are distinct features characterized by their significant width of several 1000 km/s. These lines are generated in the broad line region (BLR), a spatially extended region close to the supermassive black hole at the center of an AGN (Seyfert, 1943; Schmidt, 1963). The BLR consists of dense, rapidly moving clouds of gas that are photoionized by the intense optical-UV continuum radiation emitted from the accretion disk surrounding the black hole. As a result, these clouds emit spectral lines that are Doppler-broadened due to their high velocities, and sometimes even relativistically-broadened if close enough to the central black hole, leading to the observed broad emission line profiles in the spectra of AGNs (Baldwin, 1997; Krolik, 1999; Sulentic et al., 2000; Bianchi et al., 2019).
@@
Studies imply that BLR is universally present for accreting supermassive black holes, either in type I AGNs where the BLR is not obscured by the torus (Antonucci, 1993), or in type II AGNs where the scattered polarized light can reveal the broad lines (Antonucci & Miller, 1985); unless the luminosity, or correspondingly the accretion rate, is very low (Sabra et al., 2003; Genzel et al., 2010; Eckart et al., 2017). This indicates that the presence of BLR depends on the accretion rate (Czerny et al., 2004; Elitzur & Ho, 2009; Du et al., 2015a, 2018; Naddaf et al., 2021; Naddaf & Czerny, 2022). Although some sources with low accretion rates are claimed to have BLR (Bianchi et al., 2019), in most low luminosity sources a hot optically thin flow replaces the inner standard disk, the BLR becomes reduced to a relatively narrow ring (see e.g. Eracleous et al., 2009, for a review), and the emission seems to come directly from the illuminated disk surface. The details of the disappearance of the cold disk by the BLR are not well understood yet (e.g. Liu et al., 1999; Różańska & Czerny, 2000; Nicastro et al., 2003; Czerny et al., 2004; Liu & Qiao, 2022) and may depend on the presence of the magnetic field and hot plasma surrounding the nucleus. In the extreme cases like Sgr A* neither cold disk nor BLR are present .
@@
BLR are generally spatially unresolved apart from the recent interferometric measurements for 7 sources (Gravity Collaboration et al., 2018, 2021, 2020; GRAVITY Collaboration et al., 2024). However, the variability of the central source and the delayed response of BLR to variations in the continuum allows for the measurement of the characteristic distance from BLR to the central black hole. Given the constancy of the speed of light, , the distance is , where represents the time lag between the continuum and a selected BLR emission line, which is measured using the reverberation mapping (RM) technique (Blandford & McKee, 1982; Peterson, 1993, 2014). The technique involves monitoring the AGN continuum (typically measured at 5100 Å) and broad emission lines (most notably, H) over the extended periods. Combining the BLR radius with the broad emission-line velocity width yields the virial mass enclosed within the BLR, which is dominated by the black hole (Peterson et al., 1998; Kaspi et al., 2000). The validity of reverberation masses is supported by other means, for example by the consistency of reverberation-measured masses with other dynamical mass measurement methods (Davies et al., 2006; Onken et al., 2007; Hicks & Malkan, 2008).
@@
Optical reverberation mapping studies using the H line have shown a simple power-law relation of the form between the size of the BLR and the monochromatic luminosity of AGNs at 5100 Å (Kaspi et al., 2000; Peterson et al., 2004), known as radius-luminosity (R-L) relation. The is typically close to 0.5 for H emission lines. But despite the success of R-L relation, it exhibits a scatter, which has increased with the accumulation of more sources (Du et al., 2015b). The scatter has been attributed to accretion-rate dependence of the BLR structure, with highly accreting sources showing the largest departures from the nominal R-L relation (Du et al., 2015b, 2018). Corrections based on the Eddington ratio and dimensionless accretion rate have been proposed to reduce this scatter (Martínez-Aldama et al., 2019). Other efforts to reduce this scatter include adjustments based on the Eddington ratio and dimensionless accretion rate, with significant progress achieved by incorporating the relative Fe II strength, which correlates with accretion-rate intensity (Du & Wang, 2019).
@@
Despite significant efforts by many researchers to reduce the scatter in the R-L relation, the topic remains an open question, with ongoing investigations from both theoretical and observational perspectives. The scatter is not confined to the R-L relation derived from reverberation-measured H time-delays, but also extends to other emission lines such as Mg II and C IV. This ongoing challenge motivates further exploration of the underlying causes of the scatter in different emission lines.
@@
One more interesting topic is the use of BLR time delay for cosmology. The idea was outlined recently (Watson et al., 2011; Czerny et al., 2013; Martínez-Aldama et al., 2019), and some applications were made (Khadka et al., 2021; Cao et al., 2022, 2024) but they were based on parametric approach to the radius-luminosity relation. Without external calibration this approach can bring the results from the curvature of the Hubble diagram, constraining the but not the Hubble constant. However, with physically motivated model predicting the position of the BLR for a given source intrinsic (absolute) luminosity it would be also possible to measure the Hubble constant by combining the time delay observations with the observed source flux.
@@
In this paper, we approach the problem of the BLR properties from the theoretical side, without referring yet directly to observational data for specific objects. We present simulation results of geometry and dynamics of BLR using our 2.5D FRADO (Failed Radiatively Accelerated Dusty Outflow) model, which specifically accounts for the lowly-ionized part of the BLR (Naddaf et al., 2021). From our simulations, we calculate the theoretical transfer function, mapping BLR, determine a representative single-value delay and compare the results representing a distribution of the global parameters with the observational trends in the R-L relation. Therefore we structure the paper as follows. In section 2, we present the modeling of the BLR. Section 3 details the computation of emission line profiles based on our model, followed by section 4 where we address the calculation of time-delay maps and extraction of R-L relation from our simulations. In section 6 we discuss the results, and finally, we conclude the paper in section 7.
@@ 2 Modeling the BLR
In order to model the BLR we use 2.5D FRADO model, based on a physically motivated mechanism of dust-driven outflow from the disk surface in the region where the disk surface temperature is lower than the sublimation temperature. This assumption localizes the position of the BLR and cloud dynamics, thus the model, unlike the parametric models of BLR (e.g. Pancoast et al., 2011, 2012; Li et al., 2013), is not based on arbitrary parametrization of the cloud position. The model uses a non-hydro single-cloud approach to the dynamics of BLR.
@@
The FRADO model describes the formation of the BLR in an AGN. This model is based on the concept that radiation pressure acts on dust in the regions of the accretion disk where the temperature of the disk atmosphere is below the dust sublimation threshold. According to FRADO, this mechanism is responsible for the uplift of material that contributes to the emission of low-ionization lines. The model was initially proposed by Czerny & Hryniewicz (2011) and further enhanced to 2.5 dimension in (Naddaf et al., 2021). The model posits that radiative pressure on dust lifts clouds from the disk surface, preserving their angular momentum derived from the Keplerian motion of the disk. Once lifted, these clouds are more strongly illuminated by the inner disk regions. If a cloud becomes too hot, the dust evaporates, and the cloud continues its motion along a ballistic orbit. The global parameters of the model are the black hole mass, the Eddington rate, and the metallicity. Under conditions of lower black hole mass, lower Eddington ratio, and lower metallicity, the clouds form only a failed wind. Conversely, in scenarios of higher black hole mass, higher Eddington ratio, and higher metallicity, a fraction of the clouds may form an escaping wind. Early tests of the model (Czerny et al., 2015; Naddaf & Czerny, 2022, 2024) looked promising, in particular the application of the model to the broad absorption line quasars was attractive (Naddaf et al., 2023). Now we aim to test the model more systematically.
By knowing the global parameters of the source, we can determine the 3-D locations of statistically representative clouds and their velocities. The computational method of the dynamics of the single cloud is described in detail by Naddaf et al. (2021). We then combine the cloud positions and velocities with this the emissivity law for a specific emission line, applying the model to low-ionization lines such as H, Mg II, or Fe II.
@@
In order to perform a comprehensive study, we assumed a wide-dense range of black hole masses ( from 6 to 10 in solar units), and accretion rates ( from -2 to 0 in Eddington units), both binned with 0.2 dex. Such a grid of models corresponds well to the parameter range seen in quasar catalog (Shen et al., 2011). Sublimation temperature of dust is kept at 1500 K (Barvainis, 1987), and metallicity is set at 5 times solar, equivalent to dust-to-gas mass ratio of 0.025 (Baskin & Laor, 2018; Naddaf & Czerny, 2022). The adopted range of the viewing angles cover the range from 10∘ to 60∘, with the step of 5∘. Larger viewing angles are not considered since for such sources the BLR is generally shielded by the dusty torus, and objects are classified as type 2, with only narrow lines visible in the unpolarized spectra (see Krolik, 1999, for a general review).
@@ 3 LIL BLR line profiles
3.1 Method of line profile modeling
@@
Given a distribution of BLR clouds along with their velocity field and local emissivity, one can find the the shape of emission line profiles. As the LIL BLR locates at the radii larger than at least a few 100 (regardless of the black hole mass), the gravitational redshift may safely be neglected and only the Doppler effect contributes to line broadening. Previously, we assumed that the BLR simply follows a single-emissivity function weighted with the vertical position of a given cloud in the ensemble of clouds forming the BLR (Naddaf & Czerny, 2022). Now, in a more sophisticated manner, the line emissivity is computed convolving with the location-dependent flux of ionizing photons, .
@@
Cloud are also assumed to be optically-thin so that each cloud is spherically uniformly illuminated (i.e. no ”moon effect”) and it then transforms the absorbed radiation into isotropic line emission. Preliminary BLR simulation studies were done for 2.5D FRADO model by Naddaf et al. (2021); Naddaf & Czerny (2022). Here we do a systematic survey, aimed additionally at modeling the time delay. We use a universal ratio for the conversion of the local irradiating flux to line luminosity for all clouds. We assume that the line emissivity for H line roughly scales with the logarithm of the photon-flux which is a function of the cloud’s location (Gilbert & Peterson, 2003; Korista & Goad, 2004). In photoionization equilibrium, the flux of hydrogen recombination lines such as H generally scales with the number of ionization per unit volume, which itself is tied to the incident ionizing photon flux . At low , this relation is approximately linear; however, at higher photon fluxes, due to saturation effects in ionization and optical depth, the scaling becomes sub-linear (more slowly than linearly). This is consistent with expectations for ionization-bounded or partially matter-bounded regimes where increased ionizing photon flux leads to diminishing increases in line emission due to saturation and thermal balance constraints. Simulations performed with the code CLOUDY (Ferland et al., 1996; Moloney & Shull, 2014; Kriss et al., 2019; Temple et al., 2021; Chatzikos et al., 2023; Pandey et al., 2023) justify our approach revealing that a logarithmic scaling of H emissivity with provides an excellent empirical fit. The figure 1 shows the H line emissivity (flux) as a function of the photon flux , for photons cm-2 s-1 (Ashwani Pandey; private communications). The photon flux and the ionizing continuum spectral energy distribution (SED) with respect to the location of each cloud are directly computed from the 2.5 FRADO. These are then input into CLOUDY along with gas density, of cm-3, and column density, of cm-2. As can be seen, in the first approximation it implies a non-linear behavior of H with respect to photon flux, well approximated as a logarithmic trend (also see figure 2 in Gilbert & Peterson, 2003, for NGC 5548 as an analogy).

### v3

@@ Abstract
Context. Galaxies with active galactic nuclei (AGNs) exhibit broad emission lines as a key spectral feature. The shape of the emission-line profiles depends on the complex dynamics of discrete clouds within a spatially extended region known as the broad line region (BLR). The distribution of cloud positions within the BLR (or the de facto geometry of BLR) is directly linked to measurements of BLR time lags.
Aims. In this work, we convolved a large grid of physically based simulations of cloud distributions in BLR with photon-flux weighted emissivity of BLR clouds to investigate the generic shape of spectral line profiles. More importantly, we were able to extract the time-delay histograms of corresponding models to calculate the size of BLR.
Methods. Our physical model is based on the assumption that the clouds are launched by the radiation pressure acting on dust in the atmosphere of the outer disk. It has very few global parameters. The model is appropriate for the low-ionization part of the BLR, as seen in earlier model tests. It uses a non-hydrodynamical single-cloud approach to the BLR dynamics, which enables us to simulate the distribution of positions and velocities of the clouds.
Results. We found that the line profile width broadens with increasing black hole mass (or with viewing angle) and narrows with increasing accretion rate. The blue wing of the emission line profiles becomes more pronounced with increasing black hole mass and accretion rate, consistent with the formation and intensification of an outflow structure. We also found that the peak time delays offer a better representation of the observational trend and the scatter in the radius-luminosity (R-L) relation, compared to averaged delay values.
@@
Active Galaxies – Broad Line Region – time-delays – Eigenvector 1 – Broad Emission Lines1 Introduction
@@
Broad emission lines in the optical-UV spectra of active galactic nuclei (AGNs) are distinct features characterized by their significant width of several 1000 km/s. These lines are generated in the BLR, a spatially extended region near the central supermassive black hole in an AGN (Seyfert, 1943; Schmidt, 1963). The BLR consists of dense, rapidly moving clouds of gas that are photoionized by the intense optical-UV continuum radiation emitted from the accretion disk surrounding the black hole. As a result, these clouds emit spectral lines that are Doppler-broadened due to their high velocities, and they can even become relativistically broadened if they get close enough to the central black hole, leading to the observed broad emission line profiles in the spectra of AGNs (Baldwin, 1997; Krolik, 1999; Sulentic et al., 2000; Bianchi et al., 2019).
@@
The results of recent studies indicate that BLRs are universally present for accreting supermassive black holes. This is true for both type I AGNs, where the BLR is not obscured by the torus (Antonucci, 1993), and type II AGNs, where the scattered polarized light can reveal the broad lines (Antonucci & Miller, 1985). The only exception is when the luminosity or, correspondingly, the accretion rate, is very low (Sabra et al., 2003; Genzel et al., 2010; Eckart et al., 2017), indicating that the presence of BLR depends on the accretion rate (Czerny et al., 2004; Elitzur & Ho, 2009; Du et al., 2015, 2018; Naddaf et al., 2021; Naddaf & Czerny, 2022). Although some sources with low accretion rates have been reported to contain BLR (Bianchi et al., 2019), in most low-luminosity sources, a hot optically thin flow replaces the inner standard disk and the BLR gets reduced to a narrow ring (see e.g., Eracleous et al., 2009, for a review); thus, the emission seems to come directly from the illuminated disk surface. The details of the disappearance of the cold disk by the BLR remain poorly understood (e.g., Liu et al., 1999; Różańska & Czerny, 2000; Nicastro et al., 2003; Czerny et al., 2004; Liu & Qiao, 2022) and may depend on the presence of the magnetic field and hot plasma surrounding the nucleus. In extreme cases, such as that of Sgr A*, there is no cold disk or BLR present. In brief, Sgr A* is a very dim source with mass accretion rate of to and Eddington ratio of to (Aitken et al., 2000; Bower et al., 2003; Marrone et al., 2007; Yoon et al., 2020); this can be compared to the typical range of Eddington ratios in AGNs from up to a few.
@@
Overall, BLRs are spatially unresolved, except in recent interferometric measurements reported for seven different sources (Gravity Collaboration et al., 2018, 2021, 2020; GRAVITY Collaboration et al., 2024). However, the variability of the central source and the delayed response of BLR to variations in the continuum allows for the measurement of the characteristic distance from BLR to the central black hole. Given the constancy of the speed of light, , the distance is , where represents the time lag between the continuum and a selected BLR emission line, measured using the reverberation mapping (RM) technique (Blandford & McKee, 1982; Peterson, 1993, 2014). The technique involves monitoring the AGN continuum (typically measured at 5100 Å) and broad emission lines (most notably, H) over extended periods. Combining the BLR radius with the broad emission-line velocity width yields the virial mass enclosed within the BLR, which is dominated by the black hole (Peterson et al., 1998; Kaspi et al., 2000). The validity of reverberation masses is supported by other means; for example, by the consistency of reverberation-measured masses with other dynamical mass measurement methods (Davies et al., 2006; Onken et al., 2007; Hicks & Malkan, 2008).
@@
Optical reverberation mapping studies using the H line have revealed a power-law relation of between the size of the BLR and the monochromatic luminosity of AGNs at 5100 Å (Kaspi et al., 2000; Peterson et al., 2004). This is known as the R-L relation, where the value is typically close to 0.5 for H emission lines. Despite its success, the R-L relation exhibits scatter, which has increased with the accumulation of more sources (Du et al., 2015). This scatter is often linked to accretion rate-dependent changes in the BLR structure, with highly accreting sources showing the largest departures from the nominal R-L relation (Du et al., 2015, 2018). Several studies have explored the origin of this scatter (e.g., Wang et al., 2014; Czerny et al., 2015, 2017; Woo et al., 2024), however, corrections based on the Eddington ratio and dimensionless accretion rate to reduce the scatter yielded a tighter relation (Du et al., 2018; Mart´ınez-Aldama et al., 2019), followed by significant progress achieved by incorporating the relative Fe II strength, which correlates with accretion-rate intensity (Du & Wang, 2019a).
@@
Despite considerable efforts, scatter remains an open question, with investigations underway from both theoretical and observational perspectives. The scatter is not confined to the R-L relation derived from reverberation-measured H time-delays, but also extends to other emission lines such as Mg II and C IV. This motivates continued investigation of the underlying causes of the scatter in different emission lines.
@@
Another active area involves using BLR time delays for cosmology. The idea was outlined in several recent publications (Watson et al., 2011; Czerny et al., 2013; Mart´ınez-Aldama et al., 2019) and some applications have been made (Khadka et al., 2021; Cao et al., 2022, 2024); although we note that these have relied on a parametric approach to the R-L relation. Without external calibration, this approach can bring the results from the curvature of the Hubble diagram, constraining the but not the Hubble constant. However, with a physically motivated model predicting the position of the BLR for a given source intrinsic (absolute) luminosity, it would be also possible to measure the Hubble constant by combining the time delay observations with the observed source flux.
@@
In this paper, we investigate BLR properties from a theoretical perspective, without focusing on specific observational datasets.We present simulations of BLR geometry and dynamics using our 2.5D Failed Radiatively Accelerated Dusty Outflow (FRADO) model, which specifically accounts for the low-ionized part of the BLR (Naddaf et al., 2021). From our simulations, we were able to calculate theoretical transfer functions, map the BLR, and determine a representative single-value delay to compare these parameters across a distribution of global parameters with the observational trends in the R-L relation.
The structure of the paper is as follows. In Section 2, we present the modeling of the BLR. Section 3 details the computation of emission line profiles based on our model, followed by Section 4, where we address the calculation of time-delay maps and extraction of R-L relation from our simulations. In Section 6, we discuss our results. Finally, we present our conclusions in Section 7.
@@ 2 Modeling the BLR
To model the BLR, we used the 2.5D FRADO model, which is based on a physically motivated dust-driven outflow from the disk surface in the region where the disk surface temperature is lower than the sublimation temperature. This mechanism sets the BLR location and governs cloud dynamics. Unlike parametric BLR models (e.g., Pancoast et al., 2011, 2012; Li et al., 2013), it does not rely on arbitrary parametrization of cloud positions. It adopts a non-hydro, single-cloud dynamical approach to the BLR dynamics.
@@
The FRADO model describes the formation of the BLR in an AGN. This model is based on the concept that radiation pressure acts on dust in regions of the accretion disk where the temperature of the disk atmosphere is below the dust sublimation threshold. According to FRADO, this mechanism is responsible for the uplift of material that contributes to the emission of low-ionization lines (LIL). The model was initially proposed by Czerny & Hryniewicz (2011) and later extended to the 2.5 dimension by (Naddaf et al., 2021). The model posits that radiative pressure on dust lifts clouds from the disk surface, while preserving the angular momentum derived from the Keplerian motion of the disk. Once lifted, these clouds are more strongly illuminated by the inner disk regions. If a cloud becomes too hot, the dust evaporates and the cloud continues its motion along a ballistic orbit.
The global parameters of the model are the black hole mass, the Eddington rate, and the metallicity. Under the conditions of a lower black hole mass, lower Eddington ratio, and lower metallicity, the clouds can only form a failed wind. Conversely, in scenarios of higher black hole mass, higher Eddington ratio, and higher metallicity, a fraction of the clouds might form an escaping wind. Early tests of the model showed consistent results with observed trends (Czerny et al., 2015; Naddaf & Czerny, 2022, 2024; Naddaf, 2024), including applications to broad absorption line quasars (Naddaf et al., 2023). In this work, we test the model more systematically across a broad parameter space.
Given the global source parameters, we can compute the 3D locations and velocities of statistically representative BLR clouds. The method for computing single-cloud dynamics is detailed in Naddaf et al. (2021). We can then apply an emissivity law to the cloud distributions to model LILs, such as H, Mg II, or Fe II.
@@ 2.2 Simulation setup and the grid of models
To perform a comprehensive study, we assumed a wide and dense grid of black hole masses ( from 6 to 10 in solar units), and accretion rates ( from -2 to 0 in Eddington units), both sampled in 0.2 dex intervals. This grid spans the typical parameter range of quasars in the SDSS catalog (Shen et al., 2011). The dust sublimation temperature is fixed at 1500 K (Barvainis, 1987), and metallicity is set at five times solar, corresponding to a dust-to-gas mass ratio of 0.025 (Baskin & Laor, 2018; Naddaf & Czerny, 2022). The adopted range of the viewing angles covers the range from 10∘ to 60∘, with the step of 5∘. Larger inclinations (60∘) have been excluded, as the BLR becomes obscured by the dusty torus, and such sources appear as type 2 AGNs, with only narrow lines visible in direct light (see Krolik, 1999, for a general review).
@@ 3 LIL BLR line profiles
3.1 Line profile modeling
@@
Given a distribution of BLR clouds along with their velocity field and local emissivity, we can find the shape of the emission line profiles. As the LIL-emitting BLR lies at radii larger than a few 100 (regardless of the black hole mass), the gravitational redshift can be neglected and only the Doppler effect contributes to line broadening. Previously, we assumed that the BLR simply follows a single-emissivity function weighted with the vertical position of a given cloud in the ensemble of clouds forming the BLR (Naddaf & Czerny, 2022). We can now refine this approach by convolving the line emissivity with the location-dependent flux of ionizing photons, .
@@
Clouds are also assumed to be optically thin, so that each is uniformly illuminated in all directions (i.e., spherically symmetric illumination; no self-occultation or cloud shadowing, known as the moon effect; Rosborough et al. 2024). Each cloud then reprocesses the absorbed radiation into isotropic line emission. Preliminary BLR simulations using the 2.5D FRADO model were presented in Naddaf et al. (2021); Naddaf & Czerny (2022). Here, we have carried out a systematic survey aimed at also modeling the time delay. We adopted a universal conversion factor between local irradiating flux and line luminosity for all clouds. The H line emissivity was assumed to scale approximately with the logarithm of the local ionizing photon flux, which is a function of the cloud’s location (Gilbert & Peterson, 2003; Korista & Goad, 2004). At the photoionization equilibrium, the flux of hydrogen recombination lines, such as H, generally scales with the number of ionization per unit volume, which is itself tied to the incident ionizing photon flux . At low , this relation is approximately linear; however, at higher photon fluxes, due to saturation effects in ionization and optical depth, the scaling becomes sub-linear (more slowly than linearly). This is consistent with expectations in ionization-bounded or partially matter-bounded regimes. In such cases, increased ionizing photon flux leads to diminishing returns in line emission due to saturation and thermal balance effects. Photoionization simulations using CLOUDY (Ferland et al., 1996; Moloney & Shull, 2014; Kriss et al., 2019; Temple et al., 2021; Chatzikos et al., 2023; Pandey et al., 2023) confirm that a logarithmic dependence of H emissivity on provides an excellent empirical fit. Figure 1 shows the H line emissivity (flux) as a function of the photon flux , for photons cm-2 s-1 (Ashwani Pandey; private communications). The photon flux and the ionizing continuum spectral energy distribution (SED), with respect to the location of each cloud, were directly computed from the 2.5 FRADO. These results were then input into CLOUDY, along with fixed values for the gas density, of cm-3, and column density, of cm-2. The results confirm a non-linear H response to photon flux, which is well-approximated by a logarithmic trend (see also Figure 2 in Gilbert & Peterson, 2003, for NGC 5548).